The cost c(in dollars) of operating a certain concrete-cutting machine is related to the number of minutes n the machine is run by C(n)=2.6n^2-26n+640. What is the minimum cost?
Answer by jim_thompson5910(35256)

Min occurs at n = -b/(2a) = -(-26)/(2*2.6) = 5

So min occurs at n = 5, plug it in to get the min cost

C(n) = 2.6n^2-26n+640

C(5) = 2.6(5)^2-26(5)+640

C(5) = 575

The min cost is $575